OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a red marble? 3/10

Parth (parthkohli):

P(green) = 4/10 P(red) = 3/9 P(both) = 4/10 * 3/9

Parth (parthkohli):

Can you continue with the rest?

OpenStudy (anonymous):

not to be picky, but p(red) is not 3/9 the probability that the second one is red GIVEN that the first one is green is 3/9